OpenStudy (anonymous):

Which is an equation of a circle with center (-5, -7) that passes through the point (0, 0)?

OpenStudy (jdoe0001):

so you can pretty much see the (h, k) in the equation so what's the radius? well, notice that it says \(\bf \color{blue}{\text{passes through the point (0, 0)}}\) , meaning if you were to draw a straight line, from the center to THAT POINT, you'd get your radius so what's that distance? well \(\bf \text{distance between 2 points}\\ d = \sqrt{(x_2-x_1)^2 + (y_2-y_1)^2}\)

OpenStudy (jdoe0001):

what did you get for the radius?

OpenStudy (jdoe0001):

\(\bf d = \sqrt{(x_2-x_1)^2 + (y_2-y_1)^2}\\ \textit{2 points of (-5, -7) and (0, 0)}\\ d = \sqrt{(0-(-5))^2 + (0-(-7))^2}\)

OpenStudy (anonymous):

74.

OpenStudy (jdoe0001):

so yes, the radius is \(\bf \sqrt{74}\) so use that in your equation for your circle, keep in mind that \(\bf r = \sqrt{74} \implies r^2 = (\sqrt{74})^2 \implies r^2 = \sqrt{74^2} \implies r^2 = 74\)